OpenStudy (anonymous):

A radio commercial for a loan company states: “You only pay $0.29 a day for each $500 borrowed.” If you borrow $1,500 for 120 days, what amount will you repay, and what annual interest rate is the company actually charging?

OpenStudy (cruffo):

First - If they charge $0.29 a day for $500, how much will they charge per day for $1500?

OpenStudy (anonymous):

0.87?

OpenStudy (cruffo):

Correct: 3*500 = 1500 so 3*0.29 = 0.87. Next - If you get charged $0.87 per day, what is the total charge for 120 days?

OpenStudy (anonymous):

104.4

OpenStudy (cruffo):

Awsome! 0.87 * 120 = 104.4

OpenStudy (cruffo):

So the amount you repay is equal to the amount you borrowed plus the additional fee: 1500 + 104.40 = ??

OpenStudy (cruffo):

yep!

OpenStudy (anonymous):

wow LOL thanks so muvh my brain really isnt working tonight hahah

OpenStudy (cruffo):

word problems :(

OpenStudy (anonymous):

lol real talk

OpenStudy (cruffo):

The last part, I'm not to sure about: "and what annual interest rate is the company actually charging?"

OpenStudy (anonymous):

ya same. i was thinking a=p(1+rt) 1604.4=1500(1+r*(120/365)

OpenStudy (cruffo):

Hum. then solve for r? I was thinking: 0.29 = x% of 500. Then solving for x, looks like you pay 0.058% per day. Multiply by 365, to get 21.17%

OpenStudy (anonymous):

it comes out to the exact same answer:)

OpenStudy (cruffo):

Yah! cool.

OpenStudy (anonymous):

omg thanks you've been a big help! have a good night:)

OpenStudy (cruffo):

thanks, you too!